You susbscribed to 78 issues of Canadian geographic magazine at a total cost of $300.30. After receiving 49 issues, you decide to cancel your subscription. What refund should you receive
Answer #1
Total Subscribed for 78
Receive 49
Remaining (78-49 29
Issue Price
For 78 300.30
For 29 ?

=(29*300.30)/78

=$111.65

Refund of $111.65. Should be receive

Or

For 78 it charge $300.30 , then for 29 how much...

So we have to find price per magzine issue

Total price / Total magzine issue subscribed =300.30/78= 3.85 per magzine

Cost of 29 magzine = 29×3.85 =$111.65

Hence the amount of 29 magzine of $111.65 which is unsubscribe should be refunded.